MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2007/08/19
OBSERVER: Ohnishi Paul
CONDITION: cloudy, clear, cloudy, overcast, fog
COMMENT: Dark D2609(60sec),D2610(60sec) and D2608(300sec) are OK. High magnification microlensing MOA-378 is in gb8-chip7.
Compressor gas pressure: MPa
Expected Observation Time: START 19:02 END 06:40
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 9815-9867 | 53 | 157 | 9420 |
LMC | 1636-1639 | 4 | 16 | 4800 |
SMC | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 57 | 173 | 14220 (03h57m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
5 | 5 | 5 | 5 | 19 | 4 | 4 | 51 | 19 | 4 | 3 | 3 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 1 | 3 | 3 | 3 | 1 | 1 | 1 |
Field13 | Field14 | Field15 | Field16 |
2 | 1 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
18:00 | waiting | cloudy |
19:00 | observing | clear |
20:00 | observing | clear |
21:00 | observing | cloudy |
22:00 | waiting | overcast |
23:00 | waiting | overcast |
00:00 | waiting | fog |
01:00 | observing | cloudy |
02:00 | observing | cloudy |
03:00 | waiting | fog |
04:00 | waiting | fog |
05:00 | waiting | fog |
06:00 | observing | cloudy |
06:31 | waiting | stop due to twilight |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day